# CaDRE: Controllable and Diverse Generation of Safety-Critical Driving Scenarios using Real-World Trajectories

### v1

@@ Abstract
Simulation is an indispensable tool in the development and testing of autonomous vehicles (AVs), offering an efficient and safe alternative to road testing by allowing the exploration of a wide range of scenarios. Despite its advantages, a significant challenge within simulation-based testing is the generation of safety-critical scenarios, which are essential to ensure that AVs can handle rare but potentially fatal situations. This paper addresses this challenge by introducing a novel generative framework, CaDRE, which is specifically designed for generating diverse and controllable safety-critical scenarios using real-world trajectories. Our approach optimizes for both the quality and diversity of scenarios by employing a unique formulation and algorithm that integrates real-world data, domain knowledge, and black-box optimization techniques. We validate the effectiveness of our framework through extensive testing in three representative types of traffic scenarios. The results demonstrate superior performance in generating diverse and high-quality scenarios with greater sample efficiency than existing reinforcement learning and sampling-based methods.
@@ I INTRODUCTION
Simulation plays a pivotal role in the domain of autonomous driving, serving crucial functions in both training and evaluation [1, 2, 3, 4]. In contrast to the costly and time-consuming nature of on-road testing, simulation offers efficient feedback to developers, avoiding risky engagements in the physical world [5]. Furthermore, simulation enables the capability to incorporate various scenario sources, ranging from real-world logs and random perturbations to templates crafted by human experts. This versatility in scenario selection facilitates a comprehensive analysis of the performance of autonomous vehicles (AVs).
However, it is widely recognized that traffic scenarios in the real world exhibit a long-tail distribution, with normal scenarios constituting the majority and safety-critical scenarios occurring infrequently [6, 7]. Training AVs exclusively on these normal scenarios impedes the ability to generalize to critical situations, potentially leading to fatal accidents upon widespread deployment. During the development stage, evaluating AVs only in normal scenarios results in biased and incomplete assessments, as models may need to compromise slightly on performance in normal scenarios to improve robustness in safety-critical ones [8, 9, 10]. Consequently, there is an urgent need for the generation of safety-critical scenarios within simulations.
@@
There are three principal challenges in generating safety-critical scenarios. The first challenge is realism, which requires the scenarios generated to be sufficiently realistic to occur in the real world. This realism is typically interpreted as the similarity between the distributions of real-world and generated scenarios [11, 12, 13]. To achieve this, algorithms often involve either modifying pre-existing normal scenarios [14, 15] or employing generative models to approximate the distribution of real-world scenarios and adjusting the model-derived samples accordingly [16, 17]. The second challenge is diversity, which demands that the generation algorithm cover a wide spectrum of scenarios rather than focusing on a single failure instance. Previous approaches utilizing adversarial attack  [14, 18] or reinforcement learning [19] tend to identify the most severe cases but are lacking in producing a diverse set of scenarios. The final challenge is in ensuring that generated scenarios are in alignment with specific factors or guidelines that affect scene variation, a concept referred to as controllability. These guidelines are often expressed through constraints [20], temporal logic [16], or language [17, 21], all requiring meticulous model architecture design to facilitate the integration of these guidelines.
@@
In this paper, we introduce a generative framework CaDRE, which employs the Quality-Diversity (QD) formulation for the generation of safety-critical scenarios. Through a novel design, CaDRE addresses the above three challenges by integrating information from real-world data, domain knowledge, and black-box optimization and explicitly optimizing for both high-quality and diverse scenarios. Specifically, to maintain the realism of generated scenarios, we optimize the perturbations added to trajectories from real-world scenarios within defined constraints. Subsequently, we apply the QD algorithm to simultaneously explore and optimize continuous perturbation spaces efficiently. Finally, we achieve controllability by retrieving from archived scenarios according to the specific measure values defined by the user.
Figure 1: Overview of the CaDRE framework.
@@
  • We propose CaDRE, a novel QD formulation for the generation of diverse and controllable safety-critical scenarios in autonomous driving.
  • We propose an occupancy-aware restart mechanism as a general extension to the QD algorithm family, which improves the exploration efficiency of the algorithms.
  • We conduct experiments on three representative real-world traffic scenario types: unprotected cross-turn, high-speed lane-change, and U-turn. The experimental results demonstrate that CaDRE can generate diverse and high-quality scenarios, with better sample-efficiency compared to both RL- and sampling-based methods.
@@ II RELATED WORK
Safety-critical Scenario Generation. One significant component of autonomous driving simulation is the traffic model, which governs the behavior of the background vehicles, crucial to simulating real-world scenarios. TrafficSim [22] improves the generation process by using graph neural networks to extract interactions between vehicles. TrafficGen [11] proposes to generate the initial condition and sequential behavior of vehicles separately. ScenarioNet [23] further extends this framework to build a large-scale simulation platform that supports multiple open-source datasets.
@@
Unlike the aforementioned works, which aim to generate realistic scenarios, we focus on the long-tail distribution, consisting of the safety-critical scenarios, to provide efficient evaluations of the safety of AVs [24, 25, 26]. Most of the existing literature in this category focuses on adversarial generation. L2C [19], MMG [27], and CausalAF [28] generate initial conditions for open-loop scenario generation using reinforcement learning. The methods in [14, 29, 30] optimize the trajectories of actors with black-box optimization to attack the ego vehicle. KING [15] and AdvDO [31] further assume access to differential dynamics models to improve the efficiency of finding safety-critical scenarios. Since adversarial attacks sacrifice the diversity and controllability of generated scenarios, imitation learning [32], retrieval-augmented generation [33], causality [34], and evolutionary algorithms [35] have also been explored. To use language as conditions, LCTGen [17] predefined an intermediate representation to bridge the large language model (LLM) [36] and the low-level trajectory generator, and CTG++ [37] uses LLMs to generate signal temporal logic to guide the sampling process of diffusion models. In this paper, we depart from the common practice of leveraging adversarial generation methods and instead focus on improving the diversity and controllability of generated samples through our novel use and extensions of Quality-Diversity algorithms.
@@
Quality-Diversity Algorithms in Robotics. QD is a branch of optimization that finds a collection of high-performing, yet qualitatively different solutions [38, 39]. Specifically, QD optimizes an objective for each point in a measure space. Solving a QD problem in a continuous measure space requires infinite memory [39], so, in practice, the measure space is discretized into a finite set, and an archive is maintained to keep track of the best-known solutions over the finite set.
@@
Given QD’s ability to find a collection of high-performing solutions for different contexts, it is well-suited for many robotics applications. In the pioneering work of [40], a behavior-performance map is learned to enable the robot to quickly find a compensatory behavior and adapt after damage. QD has also been used on problems, such as human-robot interaction [41, 42], robot manipulation [43, 44], locomotion [45, 46], and morphology design [47, 48].
Popular QD algorithms, e.g., MAP-Elites [38] and CMA-ME [49], are predicated on evolutionary strategies to implement their underlying search policies. The goal of this search is to find a solution for a particular parameter configuration and to update the archive set, accordingly. These QD methods initiate this process from random regions in the search with no regard for the density of the local neighborhood of solutions; this can be inefficient, due to the possibility of restarting from already-known regions. To alleviate these issues, we introduce a novel occupancy-aware restart (OAR) mechanism, providing a form of guidance for improved coverage and efficiency during exploration. We assess the value of our novel OAR mechanism in terms of search efficiency and convergence through comparisons with the above QD algorithms and multi-particle exploration mechanisms used in reinforcement learning. To the best of our knowledge, we are the first to formulate QD for the challenging problem of safety-critical scenario generation in autonomous driving; our approach enables us to generate a map of diverse and high-quality scenes whose parameters vary smoothly along the dimensions defined by expressive measure spaces.
@@ III METHODOLOGY
Our method, Controllable and Diverse Generation of Safety-Critical Driving Scenarios using REal-world trajectories (CaDRE), integrates real-world data, domain knowledge, and black-box optimization techniques. As illustrated in Fig. 1, for each iteration, CaDRE maintains a grid archive of generated scenarios. First, it uses the QD algorithm to update the distribution from which the perturbations to the real-world trajectories are sampled. Then CaDRE simulates the perturbations to obtain diverse behavior measures and updates the archive according to the simulation results. Finally, we obtain an archive that contains thousands of critical scenarios, each with different behaviors according to the measure functions we defined using domain knowledge.
@@
Let $\bm{x}_{t}^{i}\in\mathbb{R}^{2},\psi_{t}^{i}\in[-\pi,\pi]$ and $v_{t}^{i}\in\mathbb{R}$ be the ground-plane coordinate, orientation, and speed of the world frame of the $i$-th vehicle agent at time $t$. The ego vehicle, with index $i=0$, is the vehicle for which we want to generate critical scenarios. We denote the state of the vehicle as $\bm{s}_{t}=\left\{\bm{x}_{t}^{i},\psi_{t}^{i},v_{t}^{i}\right\}_{i=0}^{N}$, where $N$ is the number of background agents. We define a specific traffic scenario as a sequence of these states $\mathcal{S}=\left\{\bm{s}_{t}\right\}_{t=0}^{T}$, where $T$ is a fixed time horizon. We initialize a specific scenario from a real-world dataset that contains only naturalistic driving scenarios.
Safety-Critical Perturbation. We perturb the trajectory of one background vehicle indexed by $i\in[1,\ldots,N]$ to generate safety-critical scenarios. We first recover the action sequence $\left\{\bm{a}_{t}^{i}\right\}^{T-1}_{t=0}$ from $\left\{\bm{s}_{t}^{i}\right\}^{T}_{t=0}$, assuming a kinematic bicycle model:
$\frac{d}{dt}\begin{bmatrix}x\\ y\\ \psi\\ v\end{bmatrix}=\begin{bmatrix}v\cos(\psi)\\ v\sin(\psi)\\ v\tan(\psi)/L\\ a\end{bmatrix},$ (1)
@@
where $L$ is the wheelbase. Each action consists of acceleration and steering input, i.e., $\bm{a}_{t}^{i}:=[a_{t}^{i},\delta_{t}^{i}]$. A new trajectory can be generated by 1) applying a sequence of bounded perturbations $\left\{\Delta\bm{a}_{t}^{i}\right\}^{T-1}_{t=0}$ to the recovered action sequence, and 2) unrolling the kinematics model from $\bm{s}_{0}^{i}$ using Eqn. 1. We then parameterize each safety-critical scenario with $\bm{\theta}=\left\{\Delta\bm{a}_{0}^{i},\ldots,\Delta\bm{a}_{T-1}^{i}\right\}\in\mathbb{R}^{T\times 2}$.
Black-Box Ego Policy. We follow [15, 31] and assume that the ego vehicle is reactive to nearby vehicles and tries to follow the original trajectory. The problem is formulated as a black-box optimization; CaDRE  does not require access to the reactive policy and would work with any other ego policies.
III-A Quality-Diversity Formulation for Scenario Generation
@@
Inspired by previous work [41, 39], we formulate the problem of generating a diverse set of safety-critical driving scenarios as a QD problem. First, we define an objective function $f:\mathbb{R}^{T\times 2}\rightarrow\mathbb{R}$ to quantify the safety-critical level. We further define K measure functions $m_{k}:\mathbb{R}^{T\times 2}\rightarrow\mathbb{R}$, jointly represented as $\bm{m}:\mathbb{R}^{T\times 2}\rightarrow\mathbb{R}^{K}$, which are a set of user-defined functions to quantify aspects of the scenario that we aim to diversify. We denote $\mathcal{M}=\bm{m}(\mathbb{R}^{T\times 2})\subseteq\mathbb{R}^{K}$ as the measure space formed by the range of $\bm{m}$. Because $f$ evaluates the quality of a scenario $\bm{\theta}$, the goal of the QD problem is to find, for each $s\in\mathcal{S}$, a scenario $\bm{\theta}$, such that $\bm{m}(\bm{\theta})=s$ and that $f(\bm{\theta})$ is maximized (Eqn. 2):
$\max\quad$ $f(\bm{\theta})$ (2)
s.t. $\bm{m}(\bm{\theta})=s,\;\forall s\in\mathcal{S}.$
@@
In practice, we discretize $\mathcal{M}$ into a finite number of $M$ cells and solve the simplified version of the problem:
@@
With a slight abuse of notation, we will use $f$ to denote the objective value and $\bm{m}$ to denote the values of the measure function. we also denote the archive as $M$, and we can retrieve the scenarios from the archive by $M[\bm{m}]$. With properly defined objective and measure functions, we can optimize a diverse population of safety-critical scenarios and retrieve individual scenarios in a controllable manner by asking for specific measure values $\bm{m}$. We build a lightweight traffic scenario simulator $Sim(\mathcal{S},\bm{\theta},i)$, which outputs the objective value $f$ and the measure values $\bm{m}$, given the original scenario $\mathcal{S}$, perturbation $\bm{\theta}$, and the index of the perturbed vehicle $i$.
Input: Real-world scenario $\mathcal{S}$, index of the perturbed background vehicle $i$, traffic simulator $Sim$, batch size $B$, an empty grid archive $M$
@@
for iter = 1, normal-$\ldots$, total_iter do
@@
III-B Design of Objective and Measure Functions
Objective Function. The objective function $f$ quantifies the safety-critical level, motivated by prior work on safety-critical scenario generation:
$f(\bm{\theta}):=\begin{cases}1,\quad\text{if vehicle $i$ collides with the egovehicle}\\ 0,\quad\text{if vehicle $i$ collides with background vehicles}\\ \exp(-\min_{t}d(\bm{x}_{t}^{0},\bm{x}_{t}^{i})),\quad\text{otherwise},\\ \end{cases}$ (4)
where $d(\cdot,\cdot)$ is the $l_{2}$ distance.
Measure Functions. The measure functions are essential to capture different aspects of critical scenarios. We propose three measure functions to define the diverse behavior of perturbed vehicles. These measure functions collectively enable the definition and evaluation of a wide range of safety-critical scenarios, focusing on essential factors such as perturbation efforts ($m_{1}$), urgency of response ($m_{2}$), and collision behavior ($m_{3}$). Here, $m_{1}$ measures the mean magnitude of the steering perturbation. It reflects how much the generated trajectory would deviate from the original trajectory:
$m_{1}=\frac{1}{t_{\text{impact}}}\sum_{t=0}^{t_{\text{impact}}-1}\left|\delta_{t}^{i}\right|.$ (5)
Next, $m_{2}$ measures the normalized impact time; it helps categorize scenarios based on the urgency of the response required, aiding in the development of time-critical decision-making algorithms for AVs:
$m_{2}=t_{\text{impact}}/T.$ (6)
Finally, $m_{3}$ measures the impact angle relative to the body frame of the vehicle ego. It allows for the evaluation of how well autonomous driving systems can recognize and react to threats from various directions, enhancing their ability to prevent accidents through appropriate maneuvering or braking:
@@ III-B Design of Objective and Measure Functions
where $R_{\psi_{t}^{0}}$ is the rotation matrix, $t=t_{\text{impact}}$, and $atan2$ is the 2-argument arctangent function. The $x$-axis of the body frame is pointing to the front of the vehicle, and the $y$-axis is pointing to the left. If there is no collision, we assume $t_{\text{impact}}=\arg\min_{t}d(\bm{x}_{t}^{0},\bm{x}_{t}^{i})$. Note that the time horizon for different vehicles can be different as some vehicles may appear or leave the scene at different times in real-world scenarios.
We adopt a variant of a QD algorithm, namely Covariance Matrix Adaptation MAP-Elites (CMA-ME) [49], to find both a higher quality and a wider diversity of safety-critical scenarios. The algorithm is adapted from CMA-ME (Algorithm 1). The key difference between QD algorithms such as CMA-ME and evolutionary strategies such as CMA-ES is that CMA-ME employs the archiving mechanism to maintain diversity [49]. Another difference is that CMA-ME adjusts the parent ranking rules that update the sampling distribution to maximize the likelihood of archive improvement; It ranks solutions filling empty cells higher than those replacing existing ones.
III-C Occupancy-Aware Restart
Existing QD algorithms [49] restart from a random elite in the archive when there is no improvement in the archive. However, it is not efficient since searching from the elites whose neighboring cells are empty is more beneficial to the exploration than from densely occupied regions in general. To improve exploration efficiency, we propose Occupancy-Aware Restart (OAR), a restart mechanism that considers the occupancy rate of neighboring cells.
As illustrated in Fig. 2, OAR assigns a higher probability to elites with more empty neighboring cells. More specifically, given the neighbor empty rate of $L$ elites $r_{1},\ldots,r_{L}$ and the temperature $T$, the softmax probability of restarting from elite $i$ is computed by:
$p_{i}=\frac{e^{r_{i}/T}}{\sum_{j=1}^{L}e^{r_{j}/T}}.$ (8)

### v2

@@ Abstract
Simulation is an indispensable tool in the development and testing of autonomous vehicles (AVs), offering an efficient and safe alternative to road testing. An outstanding challenge with simulation-based testing is the generation of safety-critical scenarios, which are essential to ensure that AVs can handle rare but potentially fatal situations. This paper addresses this challenge by introducing a novel framework, CaDRE, to generate realistic, diverse, and controllable safety-critical scenarios. Our approach optimizes for both the quality and diversity of scenarios by employing a unique formulation and algorithm that integrates real-world scenarios, domain knowledge, and black-box optimization. We validate the effectiveness of our framework through extensive testing in three representative types of traffic scenarios. The results demonstrate superior performance in generating diverse and high-quality scenarios with greater sample efficiency than existing reinforcement learning (RL) and sampling-based methods.
@@ I INTRODUCTION
Simulation plays a pivotal role in the domain of autonomous driving, serving crucial functions in both training and evaluation of autonomous vehicles (AVs) [1, 2, 3, 4]. Compared to the costly and time-consuming on-road testing, simulation offers efficient feedback to developers and avoids risky engagements in the real world [5].
It is widely recognized that traffic scenarios in the real world exhibit a long-tail distribution, where normal scenarios constitute the majority and safety-critical ones occur infrequently [6, 7]. AVs predominately trained on normal scenarios may not generalize to safety-critical ones, which can lead to fatal accidents when deployed at scale. Furthermore, evaluating AVs in such a long-tail distribution of naturalistic driving scenarios is neither sample-efficient nor comprehensive. Consequently, it is essential to generate safety-critical scenarios for both simulation-based training and evaluation.
@@
There are three principal challenges in generating safety-critical scenarios. The first challenge is realism, that is, the scenarios generated should be sufficiently realistic to occur in the real world. This realism is often quantified by the similarity between real-world scenarios and generated ones [8, 9, 10]. To achieve this, existing methods either modify existing scenarios [11, 12, 13] or employ generative models to approximate the distribution of real-world scenarios and then modify the generated samples [14, 15]. The second challenge is diversity, that is, the generation algorithm should cover a wide spectrum of scenarios rather than focusing on a single type of failure. Existing methods either lack in diversity or achieve it by construction, e.g., by specifying a multi-modal generative model [6]. The final challenge is controllability, that is the scenarios are generated according to specified characteristics. These guidelines are often expressed through constraints [16], temporal logic [14], or language [15, 17], all of which require the framework to be specially designed to facilitate their integration.
@@
To tackle these challenges, we reformulate safety-critical scenario generation within the Quality-Diversity(QD) framework and introduce our method CaDRE. QD is a branch of optimization that finds a collection of high-performing yet qualitatively different solutions [18, 19]. Through a novel design, CaDRE addresses the aforementioned challenges by integrating information from real-world data, domain knowledge, and black-box optimization. To maintain realism, we parameterize scenarios as bounded perturbations to real-world trajectories. To promote diversity, we improve on an existing QD algorithm and use it to efficiently explore and optimize the continuous parameter space. To achieve controllability, we retrieve archived scenarios according to the desired measure values.
Figure 1: (a) Overview of the CaDRE framework. CaDRE utilizes a black-box optimization algorithm to explicitly optimize for high-quality and diverse safety-critical scenarios. (b) Illustration of the Occupancy-Aware Restart (OAR) mechanism, a general extension to QD algorithms to improve exploration efficiency.
@@
  • We propose CaDRE, a novel QD formulation for safety-critical scenario generation, which achieves realism, diversity, and controllability.
  • We propose an occupancy-aware restart (OAR) mechanism as a general extension to the QD algorithm family, which improves the exploration efficiency.
  • We conduct experiments on three representative real-world traffic scenarios: unprotected cross-turn, high-speed lane-change, and U-turn, which demonstrate that CaDRE generates diverse and high-quality scenarios with better sample efficiency compared to both RL- and sampling-based methods.
@@ II RELATED WORK
Safety-critical Scenario Generation. One significant component of autonomous driving simulation is the traffic model, which governs the behavior of background vehicles. Many works, e.g., TrafficSim [20], TrafficGen [8], ScenarioNet [21], focus on generating realistic scenarios in the naturalistic distribution. In contrast, we focus on the long-tail of the distribution, that is the safety-critical scenarios, to provide efficient and comprehensive evaluations of the safety of AVs [22, 23, 24].
@@
A majority of such works focuses on adversarial generation. L2C [25], MMG [26], and CausalAF [27] generate initial conditions for open-loop scenario generation using RL. The methods in [11, 28, 29, 30] optimize the trajectories of actors with black-box optimization to attack the ego vehicle. KING [12] and AdvDO [31] further assume access to differential dynamics models to improve the efficiency of finding safety-critical scenarios. However, adversarial attack-based methods do not explicitly optimize for the diversity and controllability of generated scenarios. To address this limitation, imitation learning [32], retrieval-augmented generation [33], causality [34], and evolutionary algorithms [35] have also been explored. To control the scenario generation with language, LCTGen [15] predefines an intermediate representation to bridge the large language model (LLM) [36] and the low-level trajectory generator, and CTG++ [37] uses LLMs to generate signal temporal logic to guide the sampling process of diffusion models. In this paper, we depart from the common practice of leveraging adversarial generation methods and instead focus on improving the diversity and controllability of generated samples through our novel application and extensions of Quality-Diversity algorithms.
@@
Quality-Diversity Algorithms in Robotics. QD is a branch of optimization that finds a collection of high-performing, yet qualitatively different solutions [18, 19]. Specifically, QD optimizes an objective for each point in a measure space. Solving a QD problem in a continuous measure space requires infinite memory [19]. Thus, in practice, the measure space is discretized into a finite set, and an archive is maintained to keep track of the best-known solutions over the finite set. Some examples of popular algorithms include MAP-Elites [18], CMA-ME [38], and DQD [39].
@@
Given QD’s ability to find a collection of high-performing solutions for different contexts, it is well-suited for many robotics applications. In [40], QD is used to learn a behavior-performance map, enabling the robot to quickly find a compensatory behavior and adapt after damage. QD has also been used on problems such as human-robot interaction [41, 42], robot manipulation [43, 44], locomotion [45, 46], and morphology design [47, 48]. To the best of our knowledge, we are the first to formulate QD for safety-critical scenario generation in autonomous driving; our approach enables us to generate an archive of diverse and high-quality scenes, and allows for fine-granular control by retrieving scenarios with desired characteristics as defined by the measure space.
@@ III METHODOLOGY
Our method, Controllable and Diverse Generation of Safety-Critical Driving Scenarios using REal-world trajectories (CaDRE), integrates real-world data, domain knowledge, and black-box optimization. As illustrated in Fig. 1, for each iteration, CaDRE maintains a grid archive of generated scenarios. First, it uses the QD algorithm to update the distribution from which the perturbations to the real-world trajectories are sampled. Then CaDRE simulates the perturbations to obtain diverse behavior measures and updates the archive according to the simulation results. Finally, we obtain an archive that contains thousands of critical scenarios, each with different behaviors according to the measure functions we defined using domain knowledge.
III-A Representation of Traffic Scenarios
@@
We utilize real-world trajectories and design the parameterization of traffic scenarios to retain realism. Let $\bm{x}_{t}^{i}\in\mathbb{R}^{2},\psi_{t}^{i}\in[-\pi,\pi]$ and $v_{t}^{i}\in\mathbb{R}$ be the world-frame coordinate, orientation, and speed of the $i$-th vehicle agent at time $t$. The ego vehicle, with index $i=0$, is the vehicle for which we want to generate critical scenarios. We denote the state of the vehicle as $\bm{s}_{t}=\left\{\bm{x}_{t}^{i},\psi_{t}^{i},v_{t}^{i}\right\}_{i=0}^{N}$, where $N$ is the number of background agents. We define a specific traffic scenario as a sequence of these states $\mathcal{S}=\left\{\bm{s}_{t}\right\}_{t=0}^{T}$, where $T$ is a fixed time horizon. We initialize a specific scenario from a real-world dataset that contains only naturalistic driving scenarios.
@@
Safety-Critical Perturbation. In each generated scenario, we perturb the trajectory of a single background vehicle indexed by $i\in[1,\ldots,N]$, which is selected according to criteria to be described in Section IV-A. We first recover the action sequence $\left\{\bm{a}_{t}^{i}\right\}^{T-1}_{t=0}$ from $\left\{\bm{s}_{t}^{i}\right\}^{T}_{t=0}$, assuming a kinematic bicycle model $[\dot{x},\dot{y},\dot{\psi},\dot{v}]=[v\cos(\psi),v\sin(\psi),v\tan(\psi)/L,a]$, where $L$ is the wheelbase. Each action consists of acceleration and steering input, i.e., $\bm{a}_{t}^{i}:=[a_{t}^{i},\delta_{t}^{i}]$. A new trajectory can be generated by 1) applying a sequence of bounded perturbations $\left\{\Delta\bm{a}_{t}^{i}\right\}^{T-1}_{t=0}$ to the recovered action sequence, and 2) unrolling the kinematics model from $\bm{s}_{0}^{i}$ using the bicycle model. We then parameterize each safety-critical scenario with $\bm{\theta}=\left\{\Delta\bm{a}_{0}^{i},\ldots,\Delta\bm{a}_{T-1}^{i}\right\}\in\mathbb{R}^{T\times 2}$ and denote the allowable space of bounded perturbation as $\Theta$.
III-B Quality-Diversity Formulation for Scenario Generation
@@
Inspired by previous work [41, 19], we formulate the problem of generating a diverse set of safety-critical driving scenarios as a QD problem. First, we define an objective function $f:\mathbb{R}^{T\times 2}\rightarrow\mathbb{R}$ to quantify the safety-critical level. We further define K measure functions $m_{k}:\mathbb{R}^{T\times 2}\rightarrow\mathbb{R}$ and jointly represent them as $\bm{m}:\mathbb{R}^{T\times 2}\rightarrow\mathbb{R}^{K}$, which are a set of user-defined functions to quantify aspects of the scenario that we aim to diversify. We denote $\mathcal{M}=\bm{m}(\mathbb{R}^{T\times 2})\subseteq\mathbb{R}^{K}$ as the measure space formed by the range of $\bm{m}$. Because $f$ evaluates the quality of a scenario $\bm{\theta}$, the goal of the QD problem is to find, for each $m\in\mathcal{M}$, a scenario $\bm{\theta}$, such that $\bm{m}(\bm{\theta})=m$ and that $f(\bm{\theta})$ is maximized (Eqn. 1):
$\max$ $f(\bm{\theta})$ (1)
s.t. $\bm{m}(\bm{\theta})=m,\;\forall m\in\mathcal{M}.$
@@
In practice, we discretize $\mathcal{M}$ into a finite number of $M$ cells, which we will refer to as the archive, and solve the simplified version of the problem:
@@
With a slight overload of notation, we also use $f$ to denote the objective value and $\bm{m}$ to denote the values of the measure function. We denote the archive as $M$, and we can retrieve the scenarios from the archive by $M[\bm{m}]$. With properly defined objective and measure functions, we can optimize a diverse population of safety-critical scenarios and retrieve individual scenarios in a controllable manner by asking for specific measure values $\bm{m}$. We build a lightweight traffic scenario simulator $Sim(\mathcal{S},\bm{\theta},i)$, which outputs the objective value $f$ and the measure values $\bm{m}$, given the original scenario $\mathcal{S}$, perturbation $\bm{\theta}$, and the index of the perturbed vehicle $i$.
For this setting, the typical formulation of safety-critical scenario generation would have been $\max_{\theta\in\Theta}f(\theta)$. To promote diversity, existing methods either parameterize the scenarios as a multi-modal distribution or impose an additional distributional constraint/regularization. In comparison, the QD formulation incorporates diversity directly as part of the problem.
Input: Real-world normal scenario $\mathcal{S}$, index of the perturbed background vehicle $i$, traffic simulator $Sim$, batch size $B$, empty grid archive $M$
@@
for iter = 1, $\ldots$, total_iter do
@@
We specify the objective and measure functions here to demonstrate how concretely QD can be used to generate realistic, diverse, and controllable safety-critical scenarios. There could be many other alternative objective and measure function designs, and we hope other researchers will be inspired to adopt CaDRE for their problems.
Objective Function. The objective function $f$ quantifies the safety-critical level, motivated by prior work on safety-critical scenario generation (where $d(\cdot,\cdot)$ is the $l_{2}$ distance):
$f(\bm{\theta}):=\begin{cases}1,\quad\text{if vehicle $i$ collides with the egovehicle}\\ 0,\quad\text{if vehicle $i$ collides with background vehicles}\\ \exp(-\min_{t}d(\bm{x}_{t}^{0},\bm{x}_{t}^{i})),\quad\text{otherwise}\\ \end{cases}$ (3)
Measure Functions. The measure functions are essential to capture different aspects of critical scenarios. We propose three measure functions to define the diverse behavior of perturbed vehicles. These measure functions collectively enable the definition and evaluation of a wide range of safety-critical scenarios, focusing on representative factors such as perturbation efforts ($m_{1}$), urgency of response ($m_{2}$), and collision behavior ($m_{3}$). Here, $m_{1}$ measures the mean magnitude of the steering perturbation. It partially reflects how much the generated trajectory would deviate from the original trajectory:
$m_{1}=\frac{1}{t_{\text{impact}}}\sum_{t=0}^{t_{\text{impact}}-1}\left|\delta_{t}^{i}\right|,$ (4)
Next, $m_{2}$ measures the normalized impact time; it helps categorize scenarios based on the urgency of the response required. When there is no collision, we assume $t_{\text{impact}}=\arg\min_{t}d(\bm{x}_{t}^{0},\bm{x}_{t}^{i})$. This measure aids in the development of time-critical decision-making algorithms for AVs:
$m_{2}=t_{\text{impact}}/T,$ (5)
Finally, $m_{3}$ measures the impact angle relative to the body frame of the ego vehicle. It allows for the evaluation of how well autonomous driving systems can recognize and react to threats from various directions, enhancing their ability to prevent accidents through appropriate maneuvering or braking:
@@
where $R_{\psi_{t}^{0}}$ is the rotation matrix, $t=t_{\text{impact}}$, and $atan2$ is the 2-argument arctangent function. The $x$-axis of the body frame points to the vehicle’s front, and the $y$-axis points to the left.
III-C CaDRE Algorithm
We summarize the algorithm of CaDRE in Algorithm 1. We modify a QD algorithm, namely Covariance Matrix Adaptation MAP-Elites (CMA-ME) [38], by incorporating the domain-specific formulation (as described in Section III-A, III-B) and introducing OAR mechanism as a general extension to QD algorithms with discretized grid archives.
Algorithm Overview. The algorithm begins with the initialization of an archive and a population of solutions, which are then evaluated using $Sim$ and assigned to specific cells within the archive based on their measure values. The core of the algorithm involves iteratively selecting elite solutions according to 1) whether new cells are discovered (exploration) and 2) how much the new solutions’ objectives improve upon the old solution (exploitation), which guides the adaptation of the CMA parameters, such as the mean $\mu$ and covariance $C$. These parameters shape the distribution from which new solutions are sampled. The algorithm proceeds by replacing less optimal solutions in the archive with better-performing ones, thereby refining the solution space exploration. This process continues until a predefined total iteration is reached, ultimately resulting in an archive that captures a diverse array of scenarios across the explored measure dimensions.
Occupancy-Aware Restart. Existing QD algorithms [38] restart from random regions in the search space when there is no improvement in the archive. This is not efficient due to the possibility of restarting in well-explored regions. To improve exploration efficiency, we propose Occupancy-Aware Restart (OAR), a restart mechanism to serve as a form of guidance for improved coverage and efficiency during exploration. .